OpenStudy (goformit100):

Hameed has built a cubical water tank with lid for his house, with each outer edge 1.5 m long. He gets the outer surface of the tank excluding the base, covered with square tiles of side 25 cm. Find how much he would spend for the tiles, if the cost of the tiles is ` 360 per dozen.

OpenStudy (goformit100):

@chmvijay

OpenStudy (goformit100):

@timo86m

OpenStudy (goformit100):

@Suchandrarao

OpenStudy (nubeer):

hmm lol figure?

OpenStudy (goformit100):

No figures ... sorry wrote question wrongly

OpenStudy (goformit100):

I got the help from @Suchandrarao , Thankyou http://ncert.nic.in/NCERTS/l/iemh113.pdf 13.5 example 2

OpenStudy (goformit100):

Reply here I want to medal you

OpenStudy (anonymous):

okay yaar

OpenStudy (nurali):

The length of each edge of water tank ( a) = 1.5 m = 150 cm Given, the base of water tank is not covered with tiles, So, five outer faces will be covered by squares tiles. Therefore, Surface area of water tank = 5a^2=5(150)^2=112500cm^2 side of square tile = 25 cm So, Area of each square tile = (side)^2=25^2=625cm^2 Number of square tiles required = (Area of water tank) / (Area of each square tile =112500/625=180 Now, Given, Cost of 1 dozen ( = 12 tiles) tiles = Rs 360 so cost of 180 titles=360/12*180=5400 Hence, he will spend Rs 5400 for the tiles.